OpenStudy (ny,ny):

PLEASE HELP i need help w/ other questions too... The sum of Grover's age and that of his friend's Harold is 180 years. Thirty years ago, Grover was 3 times as old as Harold was then. How old are Grover and Harold?

OpenStudy (anonymous):

The sum (the addition of) G and H is 180 so: G+H=180 30 years ago G's age was (G-30) and H's age was (H-30). so if Grover was three times H's age 30 years ago 3(G-30)=(H-30)

OpenStudy (anonymous):

So your 2 equations with 2 unknowns are: G+H=180 3G-90 = H-30 I'd solve for H in the second equation and then substitute into the first.

OpenStudy (anonymous):

G+H=180 3G-90 = H-30 ---> 3G-90+30 = H --> 3G-60 = H G + H=180 --> G+ (3G-60) = 180 --> 4G -60 =180 --> 4G=240 -->G=60 G=60, plug into one of the 2 equations to get H

OpenStudy (ny,ny):

g=60 h=120

OpenStudy (anonymous):

"Grover was 3 times as old as Harold" Adjust a little tiny bit: G = 120 yr old , H =60 yr old

OpenStudy (ny,ny):

ohh. what if i had written that he was 60 years old on a test. do you think i'd gotten it wrong?

OpenStudy (ny,ny):

oh yeah thanks :D